Ben owns a house on a river. The river flows at 3 mph. Ben decides to start rowing downstream, with a current, at noon. He wants to return to his house at 5 P.M. At what time should Ben turn around and row home if he normally rows 5 mph in water that has no current?
Your help is appreciated.
Thanks..

Answer by stanbon(75887)

Ben owns a house on a river. The river flows at 3 mph. Ben decides to start rowing downstream, with a current, at noon. He wants to return to his house at 5 P.M. At what time should Ben turn around and row home if he normally rows 5 mph in water that has no current?
---------------------------------
Roundtrip DATA:
time = 5 hrs ; average rate = 5 mph ; distance = rt = 25 miles
--------------------------------
Downstream DATA:
distance = 12.5 miles ; rate = 8 mph ; time = d/r = 12.5/8 = 1.5625 hrs
----------------
0.5625*60 min = 33 min 45 sec
-----------------------------------
He should turn around at 1:33 + 45 sec.
============================================
Cheers,
Stan H.